OpenStudy (dls):

\[\lim_{x \rightarrow 0} \frac{\cos7x-\cos9x}{\cos3x-\cos5x}\]

OpenStudy (dls):

\[\lim_{x \rightarrow 0} \frac{\cos7x-\cos9x}{\cos3x-\cos5x}\]

OpenStudy (dls):

I did it like this: \[\lim_{x \rightarrow 0} \frac{\-\sin7x+\sin9x}{\-\sin3x+\sin5x}\]

OpenStudy (dls):

that's a negative sign

OpenStudy (dls):

@hartnn

hartnn (hartnn):

how did u get sin!! ??

OpenStudy (dls):

L Hospital

OpenStudy (anonymous):

yes L Hospitals rule

hartnn (hartnn):

divide numerator and denominator by x

hartnn (hartnn):

then use sin x/x = 1

OpenStudy (anonymous):

or u canalso use COS A-COS B formula and rewrite the entire question, then you can see if it is still 0/0 form, if so, use L hospital

OpenStudy (dls):

thanks!

OpenStudy (anonymous):

i think you are better off using COS A-COS B , formula

OpenStudy (dls):

okay,let me see

OpenStudy (dls):

well that was easy with formula!!

OpenStudy (anonymous):

glad it worked out

OpenStudy (anonymous):

math is fun : )

OpenStudy (dls):

yes!